I am having problems making this into an equation...
Kevin is 5 years older than Lisa. Four years ago, Kevin was twice as old as Lisa was then. How old is Kevin now?
My work:
Kevin Now: X+5
Lisa Now: X
Kevin Then: 2x-4
Lisa Then: X-4
Equation: 2X+5=3X-8
Coming out with wrong answer of 13; Supposed to be 14. Thanks!

Answer by venugopalramana(3286)

I am having problems making this into an equation...
Kevin is 5 years older than Lisa. Four years ago, Kevin was twice as old as Lisa was then. How old is Kevin now?
My work:
Kevin Now: X+5...GOOD
Lisa Now: X....GOOD
Kevin Then: 2x-4...NO..dont do this way..Kevin now is X+5..so 4 years back ,(as you did for Lisa)X+5-4=X+1
Lisa Then: X-4...GOOD...Twice as old Lisa then =2(X-4)=2X-8
frame the equation now and solve as you did earlier
2X-8=X+1
2X-X=1+8=9
X=9
Lisa is ( years and Kevin is X+5=9+5=14 yrs.now